Q. Choose the correct answer.
Q. lim┬(n→∞)⁡□((a^n-a^(-n))/(a^n+a^(-n) )) =?
a. 0
b. 1
c. -1
d. ∞
1
Expert's answer
2019-07-10T13:25:17-0400



"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=?"

1) If a> 1


"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{a^n(1-a^{-2n})} {a^n(1+a^{-2n})}=\lim_{n \to \infty }\frac{(1-a^{-2n})} {(1+a^{-2n})}=1"




The correct answer :b 1 


2) If 0<a< 1

"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{(1/a)^n((1/a)^{-2n}-1)} {(1/a)^n((1/a)^{-2n}+1
)}=\lim_{n \to \infty }\frac{((1/a)^{-2n}-1)} {((1/a)^{-2n}+1)}=-1"

the correct answer: c -1  .


3) If a=1

"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{1^n- 1^{-n}} {1^n+1^{-n}}=\lim_{n \to \infty }\frac{0} {2}=0"

the correct answer: a 0


4) If a= - 1

"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{(-1)^n-(-1) ^{-n}} {(-1)^n+(-1)^{-n}}=\lim_{n \to \infty }\frac{(-1)^{2n}-1} {(-1)^{2n}+1}=\lim_{n \to \infty }\frac{1-1} {1+1}=0"

5) If -1<a< 0



"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{((-1(-a))^n-(-1(-a))^{-n}} {(-1(-a))^n+(-1(-a))^{-n}}=\lim_{n \to \infty }\frac{(1/a)^n((1/a)^{-2n}-1)} {(1/a)^n((1/a)^{-2n}+1
)}"

We will introduce a replacement b=-a (0<b< 1 )



"\lim_{n \to \infty }\frac{((-1(b))^n-(-1(b)^{-n}} {(-1(b))^n+(-1(b))^{-n}}=\lim_{n \to \infty }\frac{(-1)^n(b^n-b^{-n})} {(-1)^n(b^n+b^{-n})}=\\ \\
=\lim_{n \to \infty }\frac{(b^n-b^{-n})} {(b^n+b^{-n})}=\lim_{n \to \infty }\frac{(1/b)^n((1/b)^{-2n}-1)} {(1/b)^n((1/b)^{-2n}+1
)} = -1"

the correct answer: c -1.


6) If a< - 1


If we make b= -a ,( b> 1 ) then we get a sequence


"\lim_{n \to \infty }\frac{a^n-a^{-n}} {a^n+a^{-n}}=\lim_{n \to \infty }\frac{(-1)^n b^n(1-b^{-2n})} {(-1)^n b^n(1+b^{-2n})}=\lim_{n \to \infty }\frac{(1-b^{-2n})} {(1+b^{-2n})}=1"

the correct answer :. b 1 


7) If a=0

The sequence "z_n=\frac{a^n-a^{-n}} {a^n+a^{-n}}" is undefined (division by zero).